# GUICourse: From General Vision Language Model to Versatile GUI Agent

### v1

@@ 3.3 GUIChat: A Text-Rich Multi-Modal Dataset for Interaction
We contribute this dataset in three steps: (1) Web Page Screenshot Acquisition. Utilizing Playwright to render web pages and obtain screenshots, same as the collection process of GUIEnv. (2) Textual Representation Extraction. Extracting essential structured information and text with coordinates from a redundant DOM tree. (3) Question-Answer Pair Generation. Leveraging GPT-4 to construct question-answer pairs from the textual representations of the current website. The textual representations include the elements in the website with the position information. We use GPT-4 rather than GPT-4V to achieve this step because GPT-4 can get higher accuracy than GPT-4V when given accurate textual representations.
@@ 4 Experiments
We train several GUI agents with our GUICourse (GUIEnv, GUIAct, and GUIChat datasets) based on three VLMs: Fuyu-8B [5], Qwen-VL [3], and MiniCPM-V [1]. First, we explore the performance of our GUI agents for common GUI tasks: Mind2Web [7] and AITW [6]. Then, we do ablation studies of the whole training process based on MiniCPM-V, analyzing the influence of various factors. Finally, we use two examples to show the generalization of our GUI agents.
@@ 4.1 Performance on GUI Navigation Tasks
We finetune our GUI agents (Qwen-GUI, Fuyu-GUI, and MiniCPM-GUI) and their baseline VLMs (Qwen-VL, Fuyu-8B, and MiniCPM-V) on the train set of common GUI tasks (Mind2Web and AITW) respectively and then evaluate their performance. We mainly compared our GUI agents with their baseline VLMs. Meanwhile, we add some previous works for a comprehensive comparison: We compare with MindAct [7], SeeClick [6], and GPT-4 [25] in the Mind2Web task; We compare with SeeClick, ChatGPT-CoT [48], and GPT-4V [42] in the AITW task.
Evaluation Datasets and Metrics.
Mind2Web and AITW are two commonly used GUI tasks. Mind2Web has 2000 tasks with high-level instructions in website scenarios. AITW has 30k instructions and corresponding 715k operation trajectories in smartphone scenarios. To keep consistency with previous works, we choose the processed version of these two tasks by SeeClick [6] and use the same evaluation methods as it does. We remove the "General" column of AITW in the evaluation because we convert and use a subset of the AITW-General dataset into our GUIAct dataset.
Table 2: Results of our GUI agents on the Mind2Web datasets.
Models w/o Cross-Task Cross-Website Cross-Domain
HTML Ele.Acc Op.F1 StepSR Ele.Acc Op.F1 StepSR Ele.Acc Op.F1 StepSR
MindAct 55.1 75.7 52.0 42.0 65.2 38.9 42.1 66.5 39.6
GPT-4 41.6 60.6 36.2 35.8 51.1 30.1 37.1 46.5 26.4
SeeClick 28.3 87.0 25.5 21.4 80.6 16.4 23.2 84.8 20.8
Qwen-VL 23.2 88.4 20.3 16.8 85.8 14.0 14.1 85.2 12.3
Qwen-GUI 27.9 88.1 24.4 19.3 84.1 15.6 20.5 85.4 17.5
Fuyu-8B 8.3 83.9   6.6   4.8 81.3   4.0   3.6 83.0   3.0
Fuyu-GUI 19.1 86.1 15.6 13.9 80.7 12.2 14.2 83.1 11.7
MiniCPM-V 11.0 85.6   8.5   8.2 78.2   6.0   6.5 81.4   5.2
MiniCPM-GUI 23.8 86.8 20.8 20.3 81.7 17.3 17.9 74.5 14.6
Table 3: Results of our GUI agents on the AITW datasets.
Models Modality Install Google Apps Single WebShopping Overall
ChatGPT-CoT Text   4.4 10.5   9.4   8.4   8.2
GPT-4V Image 42.6 49.8 72.8 45.7 52.7
SeeClick Image 66.4 54.9 63.5 57.6 60.6
Qwen-VL Image 70.4 57.8 70.1 64.7 65.8
Qwen-GUI Image 70.3 61.2 71.6 66.1 67.3
Fuyu-8B Image 45.9 40.0 47.2 40.8 43.5
Fuyu-GUI Image 50.9 41.6 45.7 43.8 45.5
MiniCPM-V Image 50.2 45.1 56.2 44.0 48.9
MiniCPM-GUI Image 62.3 46.5 67.3 57.5 58.4
Results.
The results in Table 2 and Table 3 demonstrate that GUICourse helps GUI agents achieve better performance than their baseline VLMs: (1) In the Mind2Web task, Qwen-GUI exhibits a remarkable increase of $2$-$5$ points in step success rate (StepSR) compared to its baseline Qwen-VL. Fuyu-8B and MiniCPM-V, which have weaker fundamental abilities, show more improvements of approximately $10$ points. The HTML-based methods (such as MindAct) achieve higher Ele.Acc and StepSR scores because they simplify the problem and use multiple-choice QA formulation. Conversely, the methods without HTML need to predict the position of the target elements based on the provided screenshots. (2) In the AITW task, our agents still have better performance than their baseline VLMs. While, these improvements are relatively smaller compared to the Mind2Web task, likely due to the fact that the samples in AITW are comparatively easier.
4.2 Ablation Study
We show two ablation experiments in this section. First, we adjust the amount of the GUIEnv-global dataset and analyze the final performances of MiniCPM-GUI. Secondly, we use the general VLM MiniCPM-V as a baseline, adding the GUIAct dataset, improving the image resolution, and mixing the GUIChat dataset successively. We analyze the influences of different datasets and factors.
Evaluation Datasets and Metrics.
We randomly sample 1.8k samples from the GUIEnv-local dataset as the test data of Bbox2Text and Text2Bbox tasks, testing the OCR and grounding ability of agents. We evaluate the Bbox2Text task by EM and F1 scores (the same as SQuAD MRC [30] task) and evaluate the Text2Bbox task by IoU scores (IoU@0.2, 0.5, and 0.7). We split about 1.4k, 1k, and 2k samples from the "web-single", "web-multi", and "smartphone" partitions in GUIAct as test data. We use the type exact match score (Type EM), step success rate (StepSR), and click accuracy (Cli. Acc) as metrics. Type EM is the accuracy of the actions’ names. StepSR is the average score of all the actions’ success rates. Cli.Acc is the average score of the success rate of the "click" and "tap" actions, which reflects the ability of grounding.
@@ 4.2 Ablation Study
Table 5: The performance of MiniCPM-GUI with different variables.
Variables Web-Single Web-Multi Smartphone Mean
+ GUIAct 89.1 55.2 52.4 62.3 48.6 42.0 64.5 43.3 29.6 41.3
+ High Resolution 91.8 69.2 72.9 68.2 47.1 42.4 72.6 52.0 47.5 54.3
+ GUIChat 91.8 70.6 74.9 67.0 47.5 45.4 71.7 53.3 44.7 55.0
@@ Results.
The performances of MiniCPM-GUI with different settings are shown in Table 4 and Table 5. We get three findings from these results:
(1) Our GUIEnv dataset is useful to improve both the fundamental abilities of VLMs and the GUI navigation ability. As shown in Table 4, the scores of the "text2bbox" and "bbox2text" tasks, as well as the performance on GUI navigation tasks (web-single) is rising during the increase in the amount of GUIEnv data. We can see that the improvement is huge from "without GUIEnv" to "2.5M GUIEnv", which demonstrates that OCR and grounding abilities are necessary. The amount of GUIEnv data does not influence the "Type EM" because predicting action names is position-free.
(2) High resolution is important for GUI navigation tasks. The default resolution of MiniCPM-V is $448\times 448$, and it can be improved to $1344\times 1344$ by the flexible patching and slicing strategy [45]. As shown in Table 5, there are huge improvements in the "web-single" and "smartphone" tasks after improving the image resolution. We also find that the “web-manual" task has little improvement. That may be because it has more complex instructions and environments, which only improve the resolution is not enough.
(3) Our GUIChat dataset is helpful for tasks in website scenarios, although its main goal is to improve the interaction ability of GUI agents. As shown in Table 5, we mix GUIChat data in the SFT state of GUI agents, and the "StepSR" improved by $2$-$3$ points on "wen-single" and "web-multi" tasks.
@@ 4.3 Case Study
Figure 5: An example of our GUI agent executing the same instruction with different environments. The numbers in the "<point></point>" are the related positions from 0 to 1000, and the "(500,500)" corresponds to the center point of the screenshot.
@@ 5 Conclusion
In this work, we construct GUICourse, a group of complete datasets (GUIEnv, GUIAct, and GUIChat) to help train GUI agents from general VLMs from three aspects respectively: improving the OCR and grounding abilities of VLMs, enhancing VLMs’ GUI knowledge, and improving GUI agents’ interaction ability. We train several GUI agents (Qwen-GUI, Fuyu-GUI, and MiniCPM-GUI) with different VLMs (Qwen-VL, Fuyu-8B, and MiniCPM-V). We evaluate the GUI agents on common GUI tasks (Mind2Web and AITW), conduct ablation studies on the self-partitioned test sets, and try them on simulated Android environments. Our experiments demonstrated that our datasets are helpful for better GUI navigation ability.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] Appendix C.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] Appendix C.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] https://github.com/yiye3/GUICourse.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Appendix B.1.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] The training cost of large models is too high, which makes it’s difficult to run experiments multiple times.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Appendix B.1.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] Section 2, Table 1
    2. (b)
      Did you mention the license of the assets? [No]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] https://github.com/yiye3/GUICourse.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Appendix C.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] Appendix C.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] Appendix A.3.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] Appendix A.3.
@@ A.1 Action Space
Figure 6: The actions in our action space.
Our datasets use the actions displayed in black color. The actions displayed in gray are not utilized, because their functions can be replaced by other actions in the same group. For example, we substitute the "type" and "paste" actions with the "input" action. Also, we replace the "drag" action (scroll bar dragging) with the "scroll" action. To facilitate the logging of information, we employ two-step actions (select_text and copy) rather than directly writing the texts. This approach is adopted to mitigate extensive time costs associated with crowd-sourcing annotation.
@@
During the contribution of our GUIAct-smartphone dataset, we use a subset of the AITW [31] dataset, a large smartphone navigation data with General/Install/Google-Apps/Single/WebShopping tags. We use a subset of the partition with the "General" tag. To keep the consistency of our action space, we convert their actions: (1) dual-point gesture: we split the "dual-point gesture" actions into "tap" and "swipe" by the distance of the touch point and the left point. Then, we use the touch point as the "point" parameter of the "tap" action, and use both points as the "dual-point" parameter of the "swipe" action. (2) type: we rename the "type" action as the "input" action, and reserve the "text" parameter of "type" as the "text" parameter of "input". (3) enter: we reserve the "enter" action. (4) go back/home: we reflect on these two actions to "tap" actions. We find the position of the two buttons on the bottom navigation bar as the "point" parameter of the "tap“ action. (5) task complete/impossible: we convert these two actions to "answer" actions, and the "task complete" and "task impossible" are the "text" parameters of the "answer" action.
@@
Two types of formats need to be considered for actions: parse format and position format. We need a structured text format to extract actions from GUI agents’ natural language responses, referred as to the parse format. Also, we need an absolute or related format to express the position information of the actions, referred as to the position format.
@@
Figure 7: The four different parse formats for action parsing.
@@ A.2 LLM-Auto Annotation
We use LLMs to help our contribution processing for GUIAct (web-single), GUIAct (web-multi), and GUIChat datasets.
GUIAct (web-single): Website Selection.

### v2

@@ 3.3 GUIChat: A Text-Rich Multi-Modal QA Dataset for Interaction
We contribute this dataset in three steps: (1) Web Page Screenshot Acquisition. Utilizing Playwright to render web pages and obtain screenshots, same as the collection process of GUIEnv. (2) Textual Representation Extraction. Extracting essential structured information and text with coordinates from a redundant DOM tree. (3) Question-Answer Pair Generation. Leveraging GPT-4 to construct question-answer pairs from the textual representations of the current website. The textual representations include the elements in the website with the position information.
@@ 4 Experiments
We train several GUI agents depending on our training pipeline with our GUICourse (GUIEnv, GUIAct, and GUIChat datasets). First, we explore the performance of our GUI agents on our test datasets and common GUI tasks. Then, we do ablation studies of the whole training process based on MiniCPM-V, analyzing the influence of various factors. Finally, we use two examples to show the generalization of our GUI agents.
Agents Web-Single Web-Multi Smartphone Mean
GPT-4o-mini* 81.0 62.3 57.0 22.0 10.0 17.0 53.0 13.0 22.0 32.0
MiniCPM-GUI w/ GUIAct 89.1 52.4 55.2 62.3 42.0 48.6 64.5 29.6 43.3 49.0
+ High Resolution 91.8 72.9 69.2 68.2 42.4 47.1 72.6 47.5 52.0 56.1
+ GUIChat 91.8 74.9 70.6 67.0 45.4 47.5 71.7 44.7 53.3 57.1
Fuyu-GUI 90.1 67.1 63.5 68.8 51.2 47.1 72.1 29.1 40.4 50.4
Qwen-GUI 90.9 69.4 66.7 68.9 52.5 46.8 73.0 55.7 58.1 57.2
Table 3: The performance of our GUI agents on test datasets. GPT-4o-mini tested on 100 samples for each dataset.
4.1 Training Settings
Depending on our training pipeline, we first improve the VLMs’ fundamental abilities (OCR and grounding ability) with the GUIEnv dataset. Then, we improve the VLMs’ GUI knowledge and conversation ability using the GUIAct and GUIChat datasets. Considering the data quantity and format of GUICourse, we designate the GUIEnv-global data for pre-training, while the remaining datasets serve as supervised fine-tuning (SFT) data.
We train the MiniCPM-GUI agent based on MiniCPM-V (Yao et al., 2024) through the following steps: (1) We pre-train MiniCPM-V by integrating 10M samples from the GUIEnv-global dataset into the original pre-training data. (2) Subsequently, we continue the training in the SFT stage by merging approximately 70k samples from the GUIEnv-local dataset, along with all samples from the GUIAct and GUIChat datasets. We also train two versions of agents with different resolutions. The default resolution of MiniCPM-V is $448\times 448$. We also have a high-resolution version with $1344\times 1344$ by the flexible patching and slicing strategy (Ye et al., 2023).
Additionally, we train other agents, Qwen-GUI and Fuyu-GUI, based on Qwen-VL Bai et al. (2023b) and Fuyu-8B Bavishi et al. (2023), respectively. However, unlike the training of MiniCPM-GUI, we only train them with the SFT data.
@@ 4.2 Performance on GUI Navigation Tasks
We evaluate our GUI agents through two methods: First, we test the performance of our agents on the test datasets split by ourselves. We split about 1.4k, 1k, and 2k samples from the "web-single", "web-multi", and "smartphone" partitions in GUIAct as test data. Then, we test the performance of our agents on other GUI navigation tasks with different action spaces, such as Mind2Web Deng et al. (2023) and AITW Rawles et al. (2023). We finetune our GUI agents and their baseline VLMs on the new GUI tasks. If our agents perform better, it demonstrates that they have learned more common GUI knowledge than baseline VLMs.
Results on our test datasets.
We use the type exact match score (Type EM), click accuracy (Cli. Acc), and step success rate (StepSR) as metrics. Type EM is the accuracy of the actions’ names. Cli.Acc is the average score of the success rate of the "click" and "tap" actions, which reflects the grounding ability. StepSR is the average score of all the actions’ success rates. More details are shown in Appendix B.2.
As shown in Table 3, we can find that: (1) Small-size models can perform effectively on visual-based GUI navigation tasks. Qwen-GUI (9.6B) achieves the highest mean step success rate of $57.2$ across the three datasets, while MiniCPM-GUI (3.1B) delivers comparable results with a mean step success rate of $57.1$. (2) High resolution is important for GUI navigation tasks. The mean step success rate improves to $56.1$ from $49.0$. (3) Our GUIChat dataset is helpful for tasks in website scenarios, although its main goal is to improve the interaction ability of GUI agents. The mean step success rate improved by $1$ point.
Agents Cross-Task Cross-Website Cross-Domain
Ele.Acc StepSR Ele.Acc StepSR Ele.Acc StepSR
SeeClick 28.3 25.5 21.4 16.4 23.2 20.8
Qwen-VL 23.2 20.3 16.8 14.0 14.1 12.3
Qwen-GUI 27.9 24.4 19.3 15.6 20.5 17.5
Fuyu-8B 8.3   6.6   4.8   4.0   3.6   3.0
Fuyu-GUI 19.1 15.6 13.9 12.2 14.2 11.7
MiniCPM-V 11.0   8.5   8.2   6.0   6.5   5.2
MiniCPM-GUI 23.8 20.8 20.3 17.3 17.9 14.6
Table 4: Results of our GUI agents on Mind2Web.
Models Install Goo.Apps Single WebShop. Overall
SeeClick 66.4 54.9 63.5 57.6 60.6
Qwen-VL 70.4 57.8 70.1 64.7 65.8
Qwen-GUI 70.3 61.2 71.6 66.1 67.3
Fuyu-8B 45.9 40.0 47.2 40.8 43.5
Fuyu-GUI 50.9 41.6 45.7 43.8 45.5
MiniCPM-V 50.2 45.1 56.2 44.0 48.9
MiniCPM-GUI 62.3 46.5 67.3 57.5 58.4
Table 5: Results of our GUI agents on the AITW.
@@
Results on other GUI navigation datasets.
Mind2Web and AITW are two commonly used GUI tasks. Mind2Web has 2000 tasks with high-level instructions in website scenarios. AITW has 30k instructions and corresponding 715k operation trajectories in smartphone scenarios. To keep consistency with previous works, we choose the processed version of these two tasks by SeeClick (Cheng et al., 2024) and use the same evaluation methods as it does. We remove the "General" column of AITW in the evaluation because we convert and use a subset of the AITW-General dataset into our GUIAct dataset.
We mainly compared our GUI agents with their baseline VLMs. The main metric is step success rate (StepSR) and we also use element accuracy (Ele.Acc) for Mind2Web, a metric to evaluate whether the action’s position parameters are correct.. The results in Table 4 and Table 5 demonstrate that GUICourse helps GUI agents achieve better performance than their baseline VLMs. (1) In the Mind2Web task, Qwen-GUI exhibits a remarkable increase of $2$-$5$ points in step success rate compared to its baseline Qwen-VL. Fuyu-8B and MiniCPM-V, which have weaker fundamental abilities, show more improvements of approximately $10$ points. (2) In the AITW task, our agents still have better performance than their baseline VLMs. However, these improvements are relatively smaller compared to the Mind2Web task, likely because we use more data (e.g., GUIEnv and GUIChat) in the website scenario.
4.3 Ablation Study
In this section, we show the influence of VLMs’ OCR and grounding abilities on simple GUI navigation tasks. We adjust the amount of the GUIEnv-global dataset in the pertaining data and analyze the final performances of MiniCPM-GUI.
Metrics of OCR and grounding abilities.
We randomly sample 1.8k samples from the GUIEnv-local dataset as the test data of Bbox2Text and Text2Bbox tasks, testing the OCR and grounding ability of agents. We evaluate the Bbox2Text task by EM and F1 scores, which is the same as the SQuAD MRC (Rajpurkar et al., 2016) task. We evaluate the Text2Bbox task by IoU scores (IoU@0.2, 0.5, and 0.7).
@@ Results.
The performances of MiniCPM-GUI with different settings are shown in Table 6. We demonstrate that our GUIEnv dataset is useful for improving the OCR and grounding ability of VLMs and the GUI navigation ability. As shown in Table 6, the scores of the "text2bbox" and "bbox2text" tasks, as well as the performance on GUI navigation tasks (web-single) is rising during the increase in the amount of GUIEnv data. We can see that the improvement is huge from "without GUIEnv" to "2.5M GUIEnv", which demonstrates that OCR and grounding abilities are necessary. The amount of GUIEnv data does not influence the "Type EM" because predicting action names is position-free.
@@ 4.4 Case Study
Figure 4: An example of our GUI agent executing the same instruction with different environments. The numbers in the "<point></point>" are the related positions from 0 to 1000.
@@
4.5 Error analysis
In this section, we analyze the error cases of our MiniCPM-GUI on one of our testing tasks (web-multi). We selected 50 error samples and manually reviewed them. After analysis, we categorized the errors into two main types: (1) False Errors. The model’s output is reasonable but does not match the predefined golden answer. (2) True Errors. The model’s output is unreasonable, including Action Type Errors (such as "click" and "swipe") and Action Position Errors (the grounding boxes).
Out of the 50 error samples analyzed, 13 samples were classified as false errors, indicating that the behavior was reasonable. There are 22 error samples related to action type errors, and 15 error samples are action position errors. We have identified some key insights from the error analysis:
(1) For the same instruction and initial screenshot, there can be multiple valid approaches to achieve the task. Static evaluation has a drawback in that some reasonable actions may be incorrectly judged as errors, referred to as false errors.
(2) Our GUI agents struggle to complete instructions when the screenshot content is complex, which can’t give a correct action, referred to as true errors.
@@ 5 Conclusion
In this work, we provide a comprehensive pipeline to train visual-based GUI agents from general VLMs. We construct GUICourse, a group of complete datasets (GUIEnv, GUIAct, and GUIChat) to help train GUI agents from general VLMs from three aspects respectively: improving the OCR and grounding abilities of VLMs, enhancing VLMs’ GUI knowledge, and improving GUI agents’ interaction ability. We train several GUI agents, evaluate them on GUI navigation tasks, and try them on simulated Android environments. Meanwhile, we conduct ablation studies to show the influence of OCR and grounding abilities.
Limitations
Acknowledgements
We sincerely thank all the anonymous reviewers and (S)ACs for their constructive comments and helpful suggestions. This work was supported by The National Natural Science Foundation of China (No. 62376273 and No.U2436209) and Shanghai Qi Zhi Institute Innovation Program SQZ202410.
@@ A.1 Action Space
Figure 5: The actions in our action space. Our datasets utilize the actions displayed in black. The actions shown in gray are not used, as their functions can be effectively replaced by other actions within the same group.
@@
During the contribution of our GUIAct-smartphone dataset, we use a subset of the AITW (Rawles et al., 2023) dataset, a large smartphone navigation data with General/Install/Google-Apps/Single/WebShopping tags. We use a subset of the partition with the "General" tag. To keep the consistency of our action space, we convert their actions: (1) dual-point gesture: we split the "dual-point gesture" actions into "tap" and "swipe" by the distance of the touch point and the left point. Then, we use the touch point as the "point" parameter of the "tap" action and use both points as the "dual-point" parameter of the "swipe" action.
(2) type: we rename the "type" action as the "input" action, and reserve the "text" parameter of "type" as the "text" parameter of "input".
(3) enter: we reserve the "enter" action.
(4) go back/home: we reflect on these two actions to "tap" actions. We find the position of the two buttons on the bottom navigation bar as the "point" parameter of the "tap“ action.
(5) task complete/impossible: we convert these two actions to "answer" actions, and the "task complete" and "task impossible" are the "text" parameters of the "answer" action.
@@
Two types of formats need to be considered for actions: parse format and position format. We need a structured text format to extract actions from GUI agents’ natural language responses, referred to as the parse format. Also, we need an absolute or related format to express the position information of the actions, referred to as the position format.
@@
Figure 6: The four different action parse formats.
@@ A.2 LLM-Auto Annotation
We use LLMs to help our contribution processing for GUIAct (single-step and multi-step website navigation), and GUIChat datasets.
A.2.1 GUIAct (single-step website navigation)
Website Selection.